# Progressive-Resolution Policy Distillation: Leveraging Coarse-Resolution Simulation for Time-Efficient Fine-Resolution Policy Learning

### v1

@@ IV Progressive-Resolution Policy Distillation
We propose a novel policy learning framework called Progressive-Resolution Policy Distillation (PRPD) to achieve time-efficient policy learning of fine-resolution simulations by utilizing coarse-resolution simulations effectively. An overview of the proposed PRPD framework is shown in Fig. 1. PRPD progressively improves simulation resolution and transfers policies at each stage. By transferring policies progressively through some middle resolutions and by transferring policies gradually while checking the stability of the policy updates instead of transferring policies all at once, we effectively transition these policies to simulations that closely resemble real-world settings. PRPD incorporates iterations of the following two steps: 1) Resolution Scheduling and 2) Policy Learning.
Simulator Assumption: To enable policy transfer, our framework follows some simulator assumptions. We assume that resolution-affected elements do not impact observations or actions. The policies’ objectives remain the same, with only the resolution differing, so the reward function is kept identical. This framework supports policy learning in high-resolution simulations, assuming that behavior is less accurate at coarser resolutions and becomes finer as resolution improves.
RL Scheme Assumption: PRPD assumes an actor-critic structure for the RL scheme, it requires policy $\pi$ and value function $Q$ for estimating $\alpha$. Applicable learning methods include the latest DRLs such as PPO [19] and SAC [16].
@@ IV-A1 Execution Flow
The simulation resolution $\Delta\in\mathbb{R}$ is scheduled based on the progress of policy learning by the resolution scheduler. Specifically, the resolution remains constant until the policy achieves the task, at which point it is progressively improved by $\Delta_{\mathcal{R}}$. The success rate is evaluated across all results of each iteration, with the success defined by meeting the threshold $\tau$. The simulation generator then creates the simulator $\mathcal{M}^{(n)}$ according to the scheduled resolution $\Delta$.
@@ IV-B1 Execution Flow
An agent collects samples $s,a,r$ from interactions with the simulation and adds them to the rollout buffer $\mathcal{D}^{(n)}$ corresponding to the current resolution. The agent then updates the current policy $\pi^{(n)}$ and current value function $Q^{(n)}$ given $\pi^{(n)}$ using the RL update scheme. At the end of each episode, the success rate $\tau$ of the policy is evaluated and passed to the resolution scheduler. When the resolution scheduler updates the resolution $\Delta$, the simulation environments are updated by the simulation generator. Then, the previous value function $Q^{(n-1)}$ and policy $\pi^{(n-1)}$ are copied to that of the current resolution $Q^{(n)}$, $\pi^{(n)}$.
PRPD approximately utilizes the conservative policy update scheme to stabilize policy transfer between different resolutions. This scheme was originally developed for a single environment $\mathcal{M}$ in a previous work [18]. This paper approximately utilizes this approach by assuming that a small resolution interval between environments makes these environments similar to a single environment ($\mathcal{M}^{(n)}\approx\mathcal{M}^{(n-1)}$). Therefore, it can be inferred that the lower the similarity (the larger the change in resolution), the lower the effectiveness, which will be verified in detail in Sec. VI-C. For this purpose, we extend the policy linear combination in Eq. (1). Specifically, whereas Eq. (1) conservatively transfers the greedy policy $\mathcal{G}(Q_{k})$, PRPD conservatively transfers the learned previous resolution policy $\pi^{(n-1)}$ as following loss function:
@@
where $\theta^{(n)}$ represents the policy network parameters for the current resolution, and KL denotes the Kullback Leibler divergence.
As shown in Eq. (2), the maximum reward and stationary distribution $d_{\pi,\mu}$ are theoretically necessary for estimating $\alpha$, but obtaining these parameters in a realistic task is seldom possible. Therefore, we modify Eq. (2) to be able to infer by mini-batch of deep RL by following previous works [18] as:
where $\alpha_{0}$ is a scaling coefficient, $\mathcal{D}^{(n)}$ is the rollout buffer of the current resolution, and $k$ is an update number. Estimated $\alpha_{k+1}$ is utilized by clipping to $[0,1]$.
Finally, PRPD leverages the dataset $D^{(n)}$ to train $Q^{(n)}$, parameterized by $\psi^{(n)}$, using the least squares approach typically employed in standard on-policy RL schemes [19]. The current policy $\pi^{(n)}$ is optimized to maximize rewards through the standard on-policy RL framework, utilizing the PPO loss function [19] alongside progressive policy transfer. Overall, the policy network $\theta^{(n)}$ and value network $\psi^{(n)}$ are updated by the summation of the these loss functions.
@@ V-B Sensors and Observations
We design three key observations for the excavation task: rock position estimated by the camera, bucket force measured by a force sensor, and bucket position and incline degree recorded by the absolute encoder. Rock position is derived from RGB images and the robot’s relative position by estimating the image-based center of gravity. The bucket force observation is represented as a binary value: only the vertical force measured by the sensor is used, with a value of 1 if it exceeds a threshold (indicating contact with a rock or the ground) and 0 otherwise (indicating the bucket is in the air). The bucket’s incline is limited to the scooping direction to reduce action dimensions and, if the torque limit is exceeded, the arm’s movement in that direction is restricted. Each observation is recorded once at the end of the agent’s action and is used as the next agent’s observation.
@@ V-C Learning Formulation
TABLE I: Calculation time per one-step sample in each resolution.
Res. $\Delta$ [mm] 70 60 50 40 30 20 10
Time [ms] 0.24 0.26 0.34 0.41 0.94 2.97 5.03
@@ V-D Variable Resolution Simulator
We designed and implemented a variable-resolution simulator for rock excavation tasks using Isaac Gym. The simulator is presented in Fig. 3 and enables us to remove the various rocks on various ground types by inclining and moving the bucket. The excavator’s whole body model and the process of moving the excavator to the rocks and putting the rocks in other places are excluded. The rock position is obtained directly by the simulation property. 3D CAD software was utilized to design the bucket. The shape of the ground soil particle was assumed to be a box. The bucket was designed to make contact only with the rock, eliminating resistance to particles representing soil; this is for the purpose of expressing that the fork-shaped bucket can move and dig out the soil.
@@
Our rock-excavation simulator generates environments containing soil and rocks based on the specified simulation resolution $\Delta$. In this paper, we focus on spatial resolution, dynamically adjusting parameters such as soil particle size and count, and rock shape precision. Rock resolution has minimal impact on simulation time but is included in the resolution to progressively increase task difficulty, as higher resolutions represent more complex shapes. For soil, particles are generated as boxes with side lengths represented by resolution $\Delta$ to match a predefined total soil volume. Rocks are represented as a collection of connected boxes. Specifically, boxes with side lengths represented by resolution $\Delta$ are connected face-to-face until the defined rock volume is reached. To create diverse rock shapes, the faces of the connected boxes are randomly selected.
@@
The calculation times and environment snapshots at different resolutions for our developed simulator are shown in Table I and Fig. 4, respectively. The calculation time of this simulator improves as resolution $\Delta$ becomes finer due to more soil particles and complex rock shapes, which result in more contact points and higher calculation times, as well as increased GPU memory usage. Consequently, finer-resolution simulations require significant calculation time, making the collection of the vast number of learning samples needed for RL very challenging.
@@
Figure 5: Comparison of learning time (top) and policy mixture rate (bottom): (Top) This compares learning time between fixed resolution learning and PRPD. $\Delta=70,\cdots,\Delta=10$ refers to resolutions as Table I, while “Mix” refers to mixed resolutions (simultaneous policy learning), both up to $400\times 128\times 128$ samples (circle points are plotted every $100\times 128\times 128$ samples). The success rate is evaluated 100 times in $\Delta=10$ at each iteration. (Bottom) This shows the task success rate and mixture rate transitions of PRPD in each resolution. The dashed line indicates when the scheduler changes resolution by achieving the target success rate $\hat{\tau}$. $\Delta=70$ is the initial environment and lacks a previous policy, so the mixture rate is omitted. Each curve plots the mean (and variance) over five experiments.
@@
VI Experiments
We conducted experiments to validate the following objectives. PRPD can learn policies in a shorter learning time compared to previous works (Sec. VI-B). The conservative policy transfer scheme stabilizes learning in PRPD (Sec. VI-C). Differences in resolution scheduling affect policy learning (Sec. VI-D). Trade-off between time efficiency and sample efficiency compared to other RL methods (Sec. VI-E). The policy trained in the simulator can be applied to various real-world environments (Sec. VI-F).
VI-A Common Settings
The experiments used a Universal Robots UR5e manipulator to control rocks with a bucket. Rock position was estimated using an Intel RealSense Depth Camera D435 and bucket force was measured with a Robotiq FT 300-S Force Torque Sensor. The robot controlled the fingertip pose at approximately 100 Hz. In all experiments, the learning parameter settings follow those of previous works, including learning parameters [19] and network architectures [20]. Detailed parameter settings are available on the project pages: https://yuki-kadokawa.github.io/prpd/. Policy learning utilized PPO with parallel environments, where the number of parallel environments was determined by the maximum that could be calculated on the PC setup (Intel Core i9-9900X CPU and GeForce RTX 3090 GPU) for the highest resolution environment ($\Delta=10$). For real-world evaluation, nine rocks of varying shapes and sizes were used, as shown in Table II.
This paper applies a domain randomization (DR) technique [20] to learn robust policies for reality gaps between the simulation and the real world. The learned policy is then utilized in real-world environments without additional learning costs. We randomized simulation parameters during training as described in the project pages. These parameters were uniformly randomized at every episode.
This experiment evaluates three key aspects. Sample number is defined as the total number of samples collected by the policy. Success rate is defined as the percentage of episodes per iteration that achieve the task, with success defined in Sec. V-C. Learning time is defined as the total training time, including both environment interactions and policy optimization.
@@ VI-C2 Results
As shown in Fig. 6(a), $\alpha=\text{opt.}$. achieves the shortest learning time compared to fixed mixture rates. Additionally, $\alpha=1$, where previous policies $\pi^{(n-1)}$ are used without optimization, outperforms $\alpha=0$, which does not use previous policies. These results confirm that the conservative policy transfer scheme enables learning with fewer samples.
@@ VI-D2 Results
As shown in Fig. 6(b), when the interval is coarse, the sample efficiency is poor, requiring about 1.5 times more samples. However, making the interval finer does not always improve efficiency, as there is almost no difference in sample efficiency when $\Delta_{\mathcal{R}}=10$ or less. This indicates that there is an optimal interval range where policy transfer works effectively. As shown in Fig. 6(c), lowering the target middle success rate $\tau$ requires more samples. This suggests that policy transfer is ineffective before the policy acquires sufficient skills.
TABLE II: Sim-to-real experiment of rock excavation: This experiment evaluated nine types of rocks (five differently shaped wooden blocks and three different-sized 3D-printed artificial rocks as shown in rock images) on three types of ground (for each element, three numbers correspond to each environment, silica sand, kinetic sand, and plastic pellets in order). “Scale” indicates approximate rock size. The first-row block shows “Ours” (PRPD) and “w/o DR” (PRPD without DR). The second-row block shows PPO at a fixed resolution. Each value indicates the success number per 20 trials, with the last column showing the average rate. The 20 trials test five policies at four positions (0, 60, 120, and 180 degrees) on a 100-mm radius circumference relative to the bucket’s initial position. Task success is defined as rocks being held in the bucket and the bucket being off the ground.

### v2

@@ IV Progressive-Resolution Policy Distillation
We propose a novel policy learning framework called Progressive-Resolution Policy Distillation (PRPD) to achieve time-efficient policy learning of fine-resolution simulations by utilizing coarse-resolution simulations effectively. An overview of the proposed PRPD framework is shown in Fig. 1. PRPD progressively improves simulation resolution and transfers policies at each stage. By transferring policies progressively through some middle resolutions and by transferring policies gradually while checking the stability of the policy updates instead of transferring policies all at once, we effectively transition these policies to simulations that closely resemble real-world settings. PRPD incorporates iterations of the following two steps: 1) Resolution Scheduling and 2) Policy Learning. The following sections outline the details of the learning steps. The pseudo-code is provided in Algorithm $1$.
Simulator Assumption: To enable policy transfer, our framework follows some simulator assumptions. Policies are designed consistently across different resolutions, even though different simulators typically vary in observations and actions. We assume that resolution-affected elements do not impact observations or actions. The policies’ objectives remain the same, with only the resolution differing, so the reward function is kept identical. This framework supports policy learning in high-resolution simulations, assuming that behavior is less accurate at coarser resolutions and becomes finer as resolution improves.
RL Scheme Assumption: PRPD assumes an actor-critic structure for the RL scheme, it requires policy $\pi$ and value function $Q$ for estimating $\alpha$. Applicable learning methods include the latest DRLs PPO [42] and SAC [40].
@@ IV-A1 Execution Flow
The simulation resolution $\Delta\in\mathbb{R}$ is scheduled based on the progress of policy learning by the resolution scheduler. Specifically, the resolution remains constant until the policy achieves the task, at which point it is progressively improved by $\Delta_{\mathcal{R}}$. The success rate is evaluated across all results of each iteration, with the task achievement defined by meeting the threshold $\tau$. The simulation generator then creates the simulator $\mathcal{M}^{(n)}$ according to the scheduled resolution $\Delta$.
@@ IV-B1 Execution Flow
An agent collects samples $s,a,r$ from interactions with the simulation and adds them to the rollout buffer $\mathcal{D}^{(n)}$ corresponding to the current resolution. The agent then updates $Q^{(n)}$ and $\pi^{(n)}$ using the RL update scheme. At the end of each episode, the success rate $\tau$ of the policy is evaluated and passed to the resolution scheduler. When the resolution scheduler updates the resolution $\Delta$, the simulation environments are updated by the simulation generator. Then, the previous value function $Q^{(n-1)}$ and policy $\pi^{(n-1)}$ are copied to that of the current resolution $Q^{(n)}$, $\pi^{(n)}$.
IV-B2 Policy Update with Conservative Policy Transfer
PRPD approximately utilizes the conservative policy update scheme to stabilize policy transfer between different resolutions. This scheme was originally developed for a single environment $\mathcal{M}$ in a previous work [47]. This paper approximately utilizes this approach by assuming that a small resolution interval between environments makes these environments similar to a single environment ($\mathcal{M}^{(n)}\approx\mathcal{M}^{(n-1)}$). Therefore, it can be inferred that the lower the similarity (the larger the change in resolution), the lower the effectiveness, which will be verified in detail in Section VI-C. For this purpose, we extend the policy linear combination in Eq. (4). Specifically, whereas Eq. (4) conservatively transfers the greedy policy $\mathcal{G}(Q^{\pi})$, PRPD conservatively transfers the learned previous resolution policy $\pi^{(n-1)}$ as follows:
$\pi_{k+1}^{(n)}(s,a)\leftarrow(1-\alpha_{k+1})\pi_{k}^{(n)}(s,a)+\alpha_{k+1}\pi^{(n-1)}(s,a).$ (6)
As shown in Eq. (5), the maximum reward and stationary distribution $d_{\pi,\mu}$ are theoretically necessary for estimating $\alpha$, but obtaining these parameters in a realistic task is seldom possible. Therefore, we modify Eq. (6) to be able to infer by mini-batch of deep reinforcement learning (DRL) by following previous works [47, 48] as:
where $\alpha_{0}$ is a scaling coefficient, $\mathcal{D}^{(n)}$ is the rollout buffer of the current resolution, and $k$ is an update number. Estimated $\alpha_{k+1}$ is utilized by clipping to $[0,1]$. Finally, the loss function of conservative policy update is defined:
@@ IV-B2 Policy Update with Conservative Policy Transfer
where $\theta^{(n)}$ represents the policy network parameters for the current resolution, and KL denotes the Kullback–Leibler divergence. To estimate $\alpha$, PRPD learns $Q^{(n)}$, the value function parameterized by $\psi^{(n)}$. The value function $Q(s,a)$ is updated using a loss function based on the temporal difference error. The loss function is defined as:
$\begin{cases}\mathcal{L}_{\psi^{(n)}}^{Q^{(n)}}=&\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}_{i}^{(n)}}\big{[}\{Q^{(n)}_{k}(s,a)-\\ &r+\gamma\mathbb{E}_{a^{\prime}\sim\pi^{(n)}(s^{\prime})}[Q^{(n)}_{k}(s^{\prime},a^{\prime})]\}^{2}\big{]}.\end{cases}$ (9)
Finally, the policy network is updated by combining the PPO loss (Eq. (3)) with the KL-based loss and the Q-function loss into the following overall loss function:
$\mathcal{L}_{\theta^{(n)},\phi^{(n)},\psi^{(n)}}=L_{\theta^{(n)},\phi^{(n)}}^{\pi^{(n)},V^{(n)}}+c_{3}\mathcal{L}_{\theta^{(n)}}^{\pi^{(n)}}+c_{4}\mathcal{L}_{\psi^{(n)}}^{Q^{(n)}},$ (10)
where $c_{3}$ and $c_{4}$ are hyperparameters.
@@ V-A Task Motions
TABLE II: Calculation time per one-step sample in each resolution.
Res. $\Delta$ [mm] 70 60 50 40 30 20 10
Time [ms] 0.24 0.26 0.34 0.41 0.94 2.97 5.03
@@ V-B Sensors and Observations
We designed three key observations for the excavation task: rock position estimated by the camera, bucket force measured by a force sensor, and bucket position and incline degree recorded by the absolute encoder. Rock position is derived from RGB images and the robot’s relative position by estimating the image-based center of gravity. The bucket force observation is represented as a binary value: only the vertical force measured by the sensor is used, with a value of 1 if it exceeds a threshold (indicating contact with a rock or the ground) and 0 otherwise (indicating the bucket is in the air). The bucket’s incline is limited to the scooping direction to reduce action dimensions and, if the torque limit is exceeded, the arm’s movement in that direction is restricted. Each observation is recorded once at the end of the agent’s action and is used as next agent’s observation.
@@ V-D Variable Resolution Simulator
We designed and implemented a variable-resolution simulator for rock excavation tasks using Isaac Gym [21]. The simulator is presented in Fig. 3 and enables us to remove the various rocks on various ground types by inclining and moving the bucket. The excavator’s whole body model and the process of moving the excavator to the rocks and putting the rocks in other places are excluded. The rock position is obtained directly by the simulation property. 3D CAD software was utilized to design the bucket. The shape of the ground soil particle was assumed to be a box. The bucket was designed to make contact only with the rock, eliminating resistance to particles representing soil; this is for the purpose of expressing that the fork-shaped bucket can move and dig out the soil.
@@
Our rock-excavation simulator generates environments containing soil and rocks based on the specified simulation resolution $\Delta$. In this paper, we focus on spatial resolution, dynamically adjusting parameters such as soil particle size and count, and rock shape precision. For soil, particles are generated as boxes with side lengths represented by resolution $\Delta$ to match a predefined total soil volume. Rocks are represented as a collection of connected boxes. Specifically, boxes with side lengths represented by resolution $\Delta$ are connected face-to-face until the defined rock volume is reached. To create diverse rock shapes, the faces of the connected boxes are randomly selected.
@@
The calculation times and environment snapshots at different resolutions for our developed simulator are shown in Table II and Fig. 4, respectively. The calculation time of this simulator improves as resolution $\Delta$ becomes finer due to more soil particles and complex rock shapes, which result in more contact points and higher calculation times, as well as increased GPU memory usage. Consequently, finer-resolution simulations require significant calculation time, making the collection of the vast number of learning samples needed for RL very challenging.
TABLE III: Range of randomized parameters in the rock excavation simulator: The parameters were used for learning policies with DR and sampled from the uniform distribution. “XY” and “XYZ” denote whether only the horizontal direction or the vertical direction is included, respectively.
Parameter w/o DR min max
[…]
Error rate of rock in bucket []delimited-[][$\cdot$][ ⋅ ] 0 0.2 0.2
[…]
Bucket torque weight (XYZ) []delimited-[][$\cdot$][ ⋅ ] 1 0.8 1.2
[…]
Friction coefficient []delimited-[][$\cdot$][ ⋅ ] 1 0.8 1.2
[…]
Total soil volume $[\textrm{mm}^{3}]$ $125^{3}$ $120^{3}$ $130^{3}$
Total rock volume $[\textrm{mm}^{3}]$ $50^{3}$ $45^{3}$ $55^{3}$
VI Experiments
We conducted experiments to validate the following objectives. PRPD can learn policies in a shorter learning time compared to previous works (Section VI-B). The conservative policy transfer scheme stabilizes learning in PRPD (Section VI-C). Differences in resolution scheduling affect policy learning (Section VI-D). Trade-off between time efficiency and sample efficiency compared to other RL methods (Section VI-E). The policy trained in the simulator can be applied to various real-world environments (Section VI-F).
VI-A Common Settings
The experiments used a Universal Robots UR5e manipulator to control rocks with a bucket. Rock position was estimated using an Intel RealSense Depth Camera D435 and bucket force was measured with a Robotiq FT 300-S Force Torque Sensor. The robot controlled the fingertip pose at approximately 100 Hz. Policy learning performance was evaluated in simulation using an Intel Core i9-9900X CPU and GeForce RTX 3090 GPU. All experiments used RL parameters as shown in Table I and the network architecture from [49]. For real-world evaluation, nine types of rocks with different shapes and sizes were used as shown in Table IV. Policy learning employed PPO with parallel environments; $E$ environments are simulated in parallel at all resolutions, which is the maximum number of parallel environments achievable on the current PC setup for the highest resolution environment ($\Delta=10$).
This paper applies a domain randomization (DR) technique [49, 50] to learn robust policies for reality gaps between the simulation and the real world. The learned policy is then utilized in real-world environments without additional learning costs. We randomized simulation parameters during training as described in Table III. These parameters were uniformly randomized at every episode.
This experiment evaluates three key aspects. Sample number is defined as the total number of samples collected by the policy. Success rate is defined as the percentage of episodes per iteration that achieve the task, with success defined in Section V-C. Learning time is defined as the total training time, including both environment interactions and policy optimization.
@@ VI-A Common Settings
Figure 5: Comparison of learning time (top) and policy mixture rate (bottom): (Top) This compares learning time between fixed resolution learning and PRPD. $\Delta=70,\cdots,\Delta=10$ refers to resolutions as Table II, while “Mix” refers to mixed resolutions (simultaneous policy learning), both up to $400\times 128\times 128$ samples (circle points are plotted every $100\times 128\times 128$ samples). The success rate is evaluated 100 times in $\Delta=10$ at each iteration. (Bottom) This shows the task success rate and mixture rate transitions of PRPD in each resolution. The dashed line indicates when the scheduler changes resolution by achieving the target success rate $\hat{\tau}$. $\Delta=70$ is the initial environment and lacks a previous policy, so the mixture rate is omitted. Each curve plots the mean (and variance) over five experiments.
@@ VI-C2 Results
As shown in Fig. 6, $\alpha=\text{opt.}$. achieves the shortest learning time compared to fixed mixture rates. Additionally, $\alpha=1$, where previous policies $\pi^{(n-1)}$ are used without optimization, outperforms $\alpha=0$, which does not use previous policies. These results confirm that the conservative policy transfer scheme enables learning with fewer samples.
@@ VI-D2 Results
As shown in Fig. 6, when the interval is coarse, the sample efficiency is poor, requiring about 1.5 times more samples. However, making the interval finer does not always improve efficiency, as there is almost no difference in sample efficiency when $\Delta_{\mathcal{R}}=10$ or less. This indicates that there is an optimal interval range where policy transfer works effectively. As shown in Fig. 6, lowering the target middle success rate $\tau$ requires more samples. This suggests that policy transfer is ineffective before the policy acquires sufficient skills.
VI-E Tradeoff between Time-Efficiency and Sample-Efficiency
VI-E1 Settings
As mentioned in Section II-C, various methods have improved RL’s sample efficiency, but less focus has been placed on “time efficiency.” This section shows that the proposed method outperforms traditional sample-efficient methods in time efficiency. SAC [40] and its ensemble variant E-SAC [31] were used as sample-efficient RL algorithms and tested within the PRPD framework and a fixed-resolution framework with $\Delta=10$.
VI-E2 Results
As shown in Fig. 6, the proposed PRPD framework outperforms fixed resolution $\Delta=10$ in time efficiency across all RL methods, despite being less sample-efficient. PPO, with the highest time efficiency among RL algorithms, achieved the fastest calculation time with PRPD. Under $\Delta=10$, PPO had the longest calculation time due to its low sample efficiency. Other sample-efficient methods required less time and fewer samples than PPO, as the slow simulation time at $\Delta=10$ increased sample collection time, making sample-efficient methods more time-efficient.
VI-F Sim-to-Real Policy Transfer
This section validates the sim-to-real policy transfer across diverse real-world rock-excavation environments, including silica sand (low viscosity), kinetic sand (high viscosity), and plastic pellets (blocky shape and large grain size). Table IV shows the evaluation of the transfer policies. The experiment shows that the success rate increases as the resolution improves in the real-world environment for fixed resolution policies, as was demonstrated with the simulation results. PRPD achieves performance comparable to fine-resolution fixed policy learning, indicating no adverse effects from using coarse-resolution simulations and successful sim-to-real transfer. We also evaluated the differences in policy performance during sim-to-real transfer depending on the presence of DR parameters. Table IV shows that policy performance dropped to approximately 50 % without DR. This indicates a reality gap between the simulation and real-world environments, mitigated by applying DR. These results demonstrate that the variable-resolution rock excavation simulator and the proposed policy learning framework achieve approximately 90 % success rates for nine types of real rock environments.
TABLE IV: Sim-to-real experiment of rock excavation: This experiment evaluated nine types of rocks (five differently shaped wooden blocks and three different-sized 3D-printed artificial rocks as shown in rock images) on three types of ground (for each element, three numbers correspond to each environment, silica sand, kinetic sand, and plastic pellets in order). “Scale” indicates approximate rock size. The first-row block shows “Ours” (PRPD) and “w/o DR” (PRPD without DR). The second-row block shows PPO at a fixed resolution. Each value indicates the success number per 20 trials, with the last column showing the average rate. The 20 trials test five policies at four positions (0, 60, 120, and 180 degrees) on a 100-mm radius circumference relative to the bucket’s initial position. Task success is defined as rocks being held in the bucket and the bucket being off the ground.
ID No. 1 No. 2 No. 3 No. 4 No. 5 No. 6 No. 7 No. 8 No. 9